EVA Report – March 5th

Crew 206 EVA Report 05 Mar 2019
EVA #07

Author : Jérémy Auclair

Purpose of EVA: The EVA has two purposes :

– Maintenance of the LOAC, installation of a rain protection

– Exploration and photography of the Hab ridge

Start time: 9:15 am

End time: 11:15 am

Narrative: We were happy to discover that the ground was very dry this morning. At 09:20 am, after having done the checks around the Hab, we took Spirit and Curiosity to go to the LOAC with some equipment. We installed a very simple rain protection for the LOAC, it’s probably not as effective as I’d like, but it’s better than nothing for now. At 09:40 am we parked Spirit and Curiosity and started walking north to climb up the ridge. We took a little less than 15 minutes to arrive to the top. The view is breathtaking, even after a year. Once on top we walked south for 25 minutes, taking picture along the way. It’s a very interesting spot because there are very different landscapes that join to this point.
After the 25 minutes walk, we turned around to find a way to go back down, we walked for 10 minutes. We walked on a south oriented light slope, for maximum security. The way down was easy, the ground was dry, everything went fine. We arrived at the Hab at 11:10 am and entered the main airlock.

Destination: LOAC and Weather Station location, Hab Ridge

Coordinates: 518500/425700 (LOAC and Weather Station), 517900/4251200 (Hab Ridge Road /Sagan Street)

EVA Participants: Norbert Pouzin (GHO), Aurélien Mure (XO), Jeremy Auclair (CMD, EVA leader)

Roads and routs per MDRS Map: Cow Dung Road

Mode of travel: Driving and Walking

Greenhab Report – March 5th

Crew 206 Green Hab Report – 05-MAR-2019
GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (6 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.0°C

High: 25.4°C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 4.38 gallons

Water in Blue Tank: 245.2 gallons

Crops watered at 0830 hours (1.73 gallons) and 1815 hours (2.65 gallons):


1) 08.30 AM 18.8° C / 30 %
2) 12.00 AM 22.5° C / 24 %
3) 15.00 PM 19.4° C / 30 %
4) 18.15 PM 18.9° C / 38 %

Changes to crops: Cress is sprouting.

Today wasn’t as hot as yesterday. I turned off the fan earlier, and only opened the door until 15.00.
I continued my experiment in the greenhab this afternoon. Everything went well.

Harvest: None.

Support/supplies needed: None.

Sol Summary – March 5th

Crew 206 March 5th 2019
Sol 9

Summary Title : Sleepy Commanders

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : Yesterday night we went to the observatory to take a look at the sky and take some pictures. There was almost no clouds, and the air was pretty clear. I had an issue with my camera lens, so I only took a couple nice pictures. As usual everyone got out of bed at 07:00 am for the daily workout. It’s still as hard, even if we’re getting used to it. After that we ate breakfast and prepared for the EVA. We went to attach a rain protection to the LOAC. It’s very simple, probably not as effective as I’d wish, but it’s good enough for now. After that we went up to the Hab ridge. The view is still as impressive and beautiful, even after a year. I took some great pictures. Cerise felt sick this morning, so she didn’t go on the EVA, and took advantage of this time to finish the Mars puzzle (unfortunately there is 3 missing pieces). After lunch, 5 of us took a long nap. It didn’t help me that much, I’m feeling sleepy since then and Aurélien, our XO, feels also quite sleepy. I haven’t made some progress on TELEOP today, our extra charger is supposed to arrive tomorrow. At least I have a set-up that works, I’m gathering a part of the data I planned on gathering. After that is was a quiet day.

Look ahead plan : We are now planning all our filming sessions for the mission video, it will take some time and some work, but it should look good !

Anomalies in work : None

Weather : Sunny in the morning, Cloudy in the afternoon

Crew physical status : Fine

EVA : LOAC maintenance and Hab Ridge exploration

Reports to be filed : GreenHab, Journalist, Operation, EVA Report, EVA Request

Support Requested : None

Journalist Report – March 4th

Crew 206 – 03/04/2019

Sol 8

Author: Benjamin Auzou, Journalist


A seism hit the station’s area in the morning. This one wasn’t a big one, about 4.6 magnitude, and nobody in the Hab felt it. However it was the biggest one in the area since 1985. We are all safe and the station is intact, our mission can continue !

What an occasion to speak about InSight mission that landed on Mars in November 2018 and which is now collecting data on Mars seismic activity. This mission is the result of the collaboration of the Unites States and Europe under the same scientific goal : listen to the heart of Mars. Between the launch and the landing of InSight, our crew and our association worked with Mid-School and High-School students on popularization of space exploration and science. We visited more than 20 schools in our French region to explain them why and how we are doing science on Mars and above all to give them the will and the desire to study Science.

Today was also a special day for our crew : Aurélien’s birthday. For the occasion in our plates : a risotto with tomatoes, a blueberries cake, a mysterious drink with blueberries and cinnamon, and more baguettes. This is the first birthday of the mission, in a few days Cerise will also celebrate her 22th birthday. What a luck to spend a birthday on an other planet !

We also began the tests in real conditions of our simulation spacesuit prototype. The EVA has been planned this afternoon because of the mud outside in the morning. Norbert was the brave of the day and worn the suit for 45 minutes. This EVA had the purpose to clarify the advantages and the issues of the spacesuit in order to improve the system. This was a success, what an achievement to finally see our SpaceSuit in its natural habitat !

Crew Photos – March 4th


Here are the photos of the day, where you can see our new simulation space suit, our daily work in the greenhab, and Aurélien’s birthday, with a blueberry cake !

Norbert, Crew 206 Greenhab Officer

EVA Report – March 4th

Crew 206 EVA Report 04-MAR-2019
EVA #5

Author: Gaspard Thieulin

Purpose of EVA:

The main purpose of the EVA was to do the basic checks of our new ‘SUPAERO SpaceSuit’. We brought a prototype of a MDRS suit which worked on it for a year. On wednesday we worked inside the hab on the protocol for pre-EVA and post EVA manipulations. Everything was nominal. Now we wanted to try it just in front of the hab.

Here is what is planned :
– Ability to realize an EVA :
– Isolation from the outside environment
– Providing air to the astronaut
– Providing water to the astronaut
– Enabaling mobility (arms, back and neck) for the astronaut – Ability to communicate with the crew
– Providing a good vision through the helmet
– Providing radio communication
– Protecting the astronaut from the sun
– Deploying the sun protection easily

Once those points were checked, we had to try this second protocole :

– Ability to remove the fog from the helmet
– Try to seat in the rovers without moving
– Try to simulate drivng conditions

But first, the team “Red” had to quickly go to the LOAC to change the battery, and check the weather station orientation after this morning’s earthquake.

Finally, the EVA teams were also meant to fix the tarp of the corridor between the science dome and the Hab.

Start time: 14h19

End time: 15h26

Narrative: Today’s EVA was conditionnaly approved, initially for the morning. We had to ask for MS’ approval before going out. In the morning, Atila went outside and experienced a very muddy ground and refused the EVA for the morning. However, he decided to let us decide if we wanted to conduct the EVA in the afternoon. The crew decided to take a decision by 1400 for an EVA replanned on 1415.

Having this decision in our hand was a lot of responsibility, especially for the EVA leader. No matter what we would decide in the afternoon, we decided to review a bit our EVA plan and add the maximum of Safety to it . Only two crew members (team Red: Aurelien and myself as EVA leader) would go outside first, taking one final decison on the pursuing of the EVA and looking for a dry area to conduct the prototype tests. Only then they would go for engineering checks and LOAC maintenance. During this whole time, the two other crew members (team Blue: Benjamin and Norbert who would try the prototype suit) would wait in the Hab, wearing EVA suits to be able to save the others as quick as possible.

Only after team Red comes back to the Hab area, one member of team Blue would swap his EVA suit for the prototype one, and go outside to conduct the rest of the EVA.

EVA leader would still be able to cancel the EVA by any time, of course.

We began equipping ourselves at 13h30 to be able to go out at 14h15, if we decided to go out. At 1400 today we decided to take a shot at this EVA’s plan, considering the whole day had been sunny and rather “hot”. Mission support (Atila in this case) accepted our plan and decision.

We went out of the airlock at 14h19. My first step on the ground would be decisive, and we were extremly pleased to discover that the ground had almost completly dried out during the day. In fact, it was very easy to move on dry dust, and the rare muddy zones were visible enough not to represent any danger. Considering that, I decided to conduct the EVA as planned and we successfully conducting every check, changed the LOAC battery and drived back while team Blue was waiting in the hab fully suited. We parked Curiosity at 14h35 and team Blue began
depressurization while Norbert was wearing the prototype suit. They also brought back some fixations for the tarps in the prototype backpack’s storage.

We were then able to conduct every test in the EVA’s protocol, and gathered a lot of data thanks to it. We ended these tests at 15h12, and Norbert was already warning us about his stamina decreasing fast due to the weight of the prototype suit.
We still were able to fix some parts of the teared tarps with the material brought in the suit in a very limited time. The fixations we used are lighter than the previous one, and we didn’t have enough to attach every part of the tarp, so we plan on another fixing session soon. Finally, we came back in the airlock at 15h21 and the EVA successfully ended at 15h26.

Destination: Hab area and LOAC / weather station.

Coordinates: 518500/425700 (LOAC and Weather Station) ; 518150/4250750 (Hab)

EVA Participants: Benjamin Auzou (COMMS), Gaspard Thieulin (ENG, Eva Leader), Aurélien Mure (XO), Norbert Pouzin (GHO, prototype Suit)

Roads and routs per MDRS Map: Cow Dung Road

Mode of travel: Walking and Rover (for carrying our heavy battery)

Vehicles used : Curiosity

Operations Report – March 4th

Crew 206 Operations Report 04 – MAR – 2019

SOL: 8

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): 18,2 hours

Solar— SOC % – Turned off Gen : 100%, Turned on Gen : 74%

Diesel Reading – 70%

Propane Reading – 81%

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 440 gallons

Water differential (static tank) – 12 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 250 gallons

Static to Loft Pump used – Yes

Water Meter: 0141653,6 gallons

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not used

Hours: 53,0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Used

Hours: 77,3 hours

Beginning charge: 100%

Ending charge: 97%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 78,4 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Nothing to report.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Not used.

Staff Crew Car used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: Still issues today. We’re beginning
to investigate local devices, I’ll keep working on it with
BDM CapCom.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: We re-attached some tarps on the connector between the green hab and the science dome. It still needs some work, as the fixations we used may break fast and not all points are attached.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report. [end]

Greenhab Report – March 4th

Crew 206 Green Hab Report – 04-MAR-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (8 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.3°C

High: 36.5°C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 4.65 gallons

Water in Blue Tank: 249.55 gallons

Crops watered at 0830 hours (2 gallons) and 1830 hours (2.65 gallons):


1) 08.30 AM 24.4° C / 31 %
2) 12.00 AM 26.1° C / 22 %
3) 16.00 PM 24.8° C / 0 %
4) 18.30 PM 18.5° C / 25 %

Changes to crops: None.

This morning, we harvested some mixed green salad and radishes for Aurélien’s birthday. We also prick out the 8 radishes pots of my experiment, to allow the best radishes to grow;

In the afternoon I performed my experiment in the greenhab, placing each pot in the growtent, with UV light and specific music for 6 minutes. However, the UV lamp found in the science dome doesn’t seem to be powerfull enough, I think it won’t be good for the experiment.

Other problem we faced today : the heater thermometer is placed on the cooling fan’s metal side. When it is hot, we put on the fan, so the heater thermometer decreases a lot more than the greenhab, and its temperature reaches 15° while it’s 25° in the greenhab, so the heater turns on.
Today, both the fan and the heater were turned on, although the temperature was high (31°C). We should re-monitor the heater thermometer.

Harvest: 89 g of mixed green salad, 1 g of watercress and 35 g of radishes.

Support/supplies needed: None.

Sol Summary – March 4th

Crew 206 March 4th 2019
Sol 8

Summary Title : Executive Birthday

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : At 07:00 am everyone got out of bed easily and with energy. The 3 of us that recorded our workout scores (as planned, the 3 others will record their scores tomorrow) beat last week’s scores. The ground was too humid to go on EVA, so we had to change a few activities on the planning. The EVA was moved to the afternoon, and the TELEOP session was moved to tomorrow. In the morning we weren’t sure that we would be able to go on EVA. In the morning we worked on our expriments and prepared a nice lunch for Aurélien’s Birthday (our executive officer). Around 14:00, the afternoon EVA was approved, and Gaspard, the EVA leader, took the first step outside on a dry ground. The EVA team managed to do everything that was planned on the request. We made some bread in the afternoon. We also learned that they had been a earthquake between 10 am and 11 am, but no one noticed anything. The crew is still feeling well and motivated, but the routine has definitely settled.

Look ahead plan : Tomorrow will be a usual day, we hope to make some progress on TELEOP.

Anomalies in work : None

Weather : Partly sunny in the morning, sunny in the afternoon

Crew physical status : Fine

EVA : LOAC maintenance and simulated space suit in the afternoon

Reports to be filed : GreenHab, Journalist, Operation, EVA Report, EVA Request

Support Requested : Our Astronomer would really like some sun and a clear sky.

Journalist Report – March 03rd

Crew 206 – 03/03/2019

Sol 7

Author: Benjamin Auzou, Journalist

"Martian Sol-Off"

For our first day-off of the mission, we played cards, begin to complete the Mars puzzle of the MDRS. But above all we spend a lot of time in the kitchen: Gaspard and Aurélien prepared pancakes for breakfast, Jérémy and Cerise baked French baguettes, and Aurélien and I cooked a gratin dauphinois (potatoes cooked in cream and cheese baked in the oven). All that was bettered by the basil leafs brought by Norbert from the Green-Hab.

Outside the Hab, the weather changed all along the day. After a sunny morning and a rainy noon, the afternoon was shared between grey clouds and hail. Even if it wasn’t Mars-like, the sound of rain and hail on the Habs roof was relaxing. In the same time our furious Monopoly Deal games were the great activity of the afternoon, even if Gaspard’s luck with cards drawing killed the competition.

This day-off is for us the occasion to present you the results of our water monitoring. We are counting the water we use in different categories: drinking, cooking, dishes, hygiene, flushes and Green-Hab. Thanks to that we managed to reduce our consumption and realized a precise monitoring. We measured in average 88% of our water consumption and classified that in the categories for each sol spend in the MDRS.

Today we recharged our batteries for the week to come that will be busy between the first tests our own simulation spacesuit and the follow-up of the human factors experiments.

More than one third of our mission is done, this first week was balanced between the first marks in the station, the first scientific experiments but also the first difficulties. As in all space missions, we are facing difficulties and we are led to adapt ourselves and react with what we have here in the station. More than ever, we are feeling that we make science evolve in the right way to Mars.

Copyright © The Mars Society. All rights reserved. | Main Site